OpenStudy (anonymous):

How many grams of H2 are needed produced in 11.13 grams of of NH3

OpenStudy (aaronq):

can you write an equation for the reaction and balance it?

OpenStudy (anonymous):

3H2+N2\[3H _{2}+N _{2}\rightarrow2NH _{3}\]

OpenStudy (anonymous):

ignore the first part

OpenStudy (aaronq):

okay good stuff, now covert the grams of NH3 to moles

OpenStudy (anonymous):

ok so find the molar mass?

OpenStudy (aaronq):

well yeah, then find the moles moles=mass/molar mass

OpenStudy (anonymous):

17.034

OpenStudy (anonymous):

.653

OpenStudy (aaronq):

okay so you have this 3H2+N2→2NH3 x 0.653 now, you have to use the coefficients to form a ratio between the 2 substances, since they're not in a one to one relationship. aA + bB -> cC moles of C/c = moles of A/a

OpenStudy (anonymous):

see i dont understand that

OpenStudy (aaronq):

all you have to do is normalize each one by diving the amount of moles by it's respective coefficient \[\frac{ moles of H _{2} }{ 3 }=\frac{ moles of NH _{3} }{ 2 }\]

OpenStudy (aaronq):

then solve for whatever you're looking for